# FastLongSpeech: Enhancing Large Speech-Language Models for Efficient Long-Speech Processing

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has spurred significant progress in Large Speech-Language Models (LSLMs), enhancing their capabilities in both speech understanding and generation. While existing LSLMs often concentrate on augmenting speech generation or tackling a diverse array of short-speech tasks, the efficient processing of long-form speech remains a critical yet underexplored challenge. This gap is primarily attributed to the scarcity of long-speech training datasets and the high computational costs associated with long sequences. To address these limitations, we introduce FastLongSpeech, a novel framework designed to extend LSLM capabilities for efficient long-speech processing without necessitating dedicated long-speech training data. FastLongSpeech incorporates an iterative fusion strategy that can compress excessively long-speech sequences into manageable lengths. To adapt LSLMs for long-speech inputs, it introduces a dynamic compression training approach, which exposes the model to short-speech sequences at varying compression ratios, thereby transferring the capabilities of LSLMs to long-speech tasks. To assess the long-speech capabilities of LSLMs, we develop a long-speech understanding benchmark called LongSpeech-Eval. Experiments show that our method exhibits strong performance in both long-speech and short-speech tasks, while greatly improving inference efficiency.
@@ 1 Introduction
To handle speech inputs, traditional methods [8, 9] typically employ a cascaded pipeline, where speech is first transcribed into text and then processed by LLMs. However, these approaches suffer from error propagation and discard valuable paralinguistic information [10]. To overcome these drawbacks, recent research [11, 12, 13] has shifted towards an end-to-end paradigm, enabling LSLMs to directly process and reason with speech signals. These methods can be broadly divided into two categories. On one hand, some approaches [4, 14, 15] such as Qwen2-Audio [5] align the output spaces of pre-trained audio encoders with the embedding of LLMs, which allows speech inputs to be accommodated while transferring partial capabilities of the employed LLMs. At the same time, other methods [11, 16] involves discretizing the speech to discrete units, which allows LSLMs to handle speech units similar to text tokens. Despite these advancements, current LSLMs are largely constrained to processing short speech segments, typically under 30 seconds [5, 17]. Only a few LSLMs [12] have achieved a processing duration of 30 minutes on speech summarization tasks by relying on the construction of extensive, specialized training datasets.
@@ 2 Background
Connectionist Temporal Classification (CTC) [22] is widely used in Automatic Speech Recognition (ASR), where the input is typically longer than the corresponding output. To align speech with the transcript, CTC introduces a blank token into the vocabulary and defines the possible output as an alignment $\mathbf{a}$. Each time step in this alignment corresponds to either a blank token or a non-blank token. The alignment has the same length as the speech sequence and can be reduced to the final transcript through a collapse function $\Gamma^{-1}$. During training, CTC employs an efficient dynamic programming algorithm to maximize the probability of all possible alignments corresponding to the ground-truth transcript $\mathbf{c}$:
@@ 3.2 Iterative Fusion
$r(m)=T(m)-T(m\!+\!1).$ (6)
Subsequently, we utilize the similarity metric between adjacent frames, as defined in Eq.(4) to identify the $r(m)$ most similar pairs of adjacent frames. The consecutive identified frames are grouped into a span, as illustrated in Figure 1. For each span, we employ a weighted fusion approach, leveraging the content density in Eq.(3) as weights to merge all frames within the span into a single compressed speech frame. This process yields a new sequence of speech representations with length $T(m\!+\!1)$, comprising both the newly condensed frames and the remaining uncompressed frames. If $T(m\!+\!1)$ still exceeds the target length $L$, we initiate another iteration. Otherwise, the resulting speech representations from this round constitute the final condensed representations $\mathbf{h}^{\prime}$.
@@ 3.3 Training Method
In the second stage, we introduce a novel dynamic compression training method. The introduction of this method is based on two considerations. First, it enables the LLM to adapt to condensed representations $\mathbf{h}^{\prime}$ with different compression ratios. Second, the current <$\mathbf{s}$, $\mathbf{x}$, $\mathbf{y}$> triplet training data primarily contains short-speech clips, which are typically under 30 seconds in duration [27]. By sampling the length $L$ of condensed representations, LSLM can maintain its perception of speech sequences corresponding to the length of its speech window, without avoiding excessive bias towards overly condensed speech sequences. The dynamic compression training method is as follows:
@@ 4.3 Main Results
For the short-speech spoken QA task, Figure LABEL:main_res illustrates the performance of various methods across three datasets. Our FastLongSpeech method consistently outperforms other methods on all three datasets under different speech compression ratios, maintaining high response quality even at a 30-fold compression ratio ($L$ = 25). Unlike the Random method, which arbitrarily discards speech representations, other methods consider all temporal information when compressing speech representation, resulting in improved generation quality [27]. Compared to AvgPool and MostSim methods, our method more effectively eliminates redundant information while preserving highly informative speech representations, leading to better performance across various compression ratios. Notably, when $L$ equals 12, other speech fusion methods exhibit similarly suboptimal performance, while our method maintains a substantial performance advantage. We attribute this superiority to our novel iterative fusion strategy and dynamic compression training approach. Furthermore, compared to vanilla Qwen2-Audio, our method achieves comparable performance with a shorter sequence of speech representations, demonstrating higher efficiency.
@@ 5.2 Inference Efficiency
Ours ($L$=400) 3.80 8.54
Ours ($L$=200) 3.87 5.64
Ours ($L$=100) 3.71 4.17
@@
In short-speech tasks, our approach demonstrates performance comparable to vanilla Qwen2-Audio while requiring only half the computational resources. When the allocated computational resources are increased, we can achieve better results. Notably, the computational costs decrease as the compression ratio increases. This not only demonstrates the better efficiency of our model but also highlights its ability to balance generation quality and inference efficiency. The advantages of our method become even more pronounced in long-speech tasks, where our method achieves better generation quality than NTK-ROPE, with a 70% reduction in runtime and a 60% decrease in computational costs. This further shows the effectiveness of our method in handling long-speech inputs. For spoken dialogue understanding and emotion recognition tasks, please refer to the Appendix E and F.
@@ 5.3 Content of Condensed Representations
Beyond the spoken QA, spoken dialogue understanding and emotion recognition tasks, we extend our evaluation to the ASR task, which requires precise transcription of the entire speech content [43]. Through this task, we explore variations in condensed representations across different compression ratios. Table 5 demonstrates the ASR performance of Qwen2-Audio and our method.
@@
Ours ($L$=750) 4.04 7.02 11.76
Ours ($L$=400) 4.08 7.17 11.77
Ours ($L$=200) 4.36 7.40 12.70
Ours ($L$=100) 27.12 24.61 23.69
At low compression ratios ($L$=400), FastLongSpeech performs comparably to Qwen2-Audio, demonstrating the effectiveness of our dynamic compression training and iterative fusion strategy in preserving speech content. Unlike Qwen2-Audio, our method does not require substantial post-processing to extract the transcript, with strong instruction following abilities. At higher compression ratios ($L$=100), FastLongSpeech slightly trails Qwen2-Audio in ASR but maintains comparable results in spoken QA, as illustrated in Figure LABEL:main_res. This indicates that while our approach demonstrates applicability across diverse tasks, the optimal compression ratio is inherently task-dependent. Therefore, achieving an effective balance between efficiency and effectiveness thus necessitates careful calibration and a thorough assessment of resource constraints.
@@ 6 Related Work
Long sequence modeling presents challenges across diverse domains, including text, video, and speech. The approaches to long-context modeling vary depending on the type of the inputs. For extended text sequences, researchers explored methods such as position interpolation and extrapolation [47], sliding window [48], continuous fine-tuning on long-text data [49], and native sparse attention [50]. To address long-video processing, recent works leverage frame selection or merging strategies [51], as well as vision token merging techniques [52]. In the realm of speech processing, early methods focus on enhancing the performance of ASR [53] and speech translation [54] through speech compression techniques. More recently, FastAdaSP [37] mitigates inference overhead by performing token selection within LLMs. Concurrently, Speechprune [27] employs a token selection strategy to extend the effective speech window of Qwen2-Audio to 90 seconds for Speech Information Retrieval task.
@@ Appendix E Applicability of Our Method to Vanilla LSLMs
Ours ($L$=400) 4.13 8.25
Ours ($L$=200) 3.92 5.46
Ours ($L$=100) 3.62 4.06
Ours ($L$=50) 3.16 3.35
To demonstrate the effectiveness and robustness of our method, we first extend our experiments to spoken dialogue understanding task. For this task, we conduct experiments on vanilla Qwen2-Audio using speech_dialogue_QA_fisher of AIR-Bench [35]. As shown in Table 7, our method effectively balances performance and inference efficiency. Notably, at lower compression ratios ($L$ = 200), our approach demonstrate comparable performance to the vanilla Qwen2-Audio model with a 50% reduction in computational costs. Moreover, even at a higher compression ratio of 15x ($L$=50), our method still maintains robust performance. These findings underscore the efficacy and versatility of our iterative fusion strategy.

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has spurred significant progress in Large Speech-Language Models (LSLMs), enhancing their capabilities in both speech understanding and generation. While existing LSLMs often concentrate on augmenting speech generation or tackling a diverse array of short-speech tasks, the efficient processing of long-form speech remains a critical yet underexplored challenge. This gap is primarily attributed to the scarcity of long-speech training datasets and the high computational costs associated with long sequences. To address these limitations, we introduce FastLongSpeech, a novel framework designed to extend LSLM capabilities for efficient long-speech processing without necessitating dedicated long-speech training data. FastLongSpeech incorporates an iterative fusion strategy that can compress excessively long-speech sequences into manageable lengths. To adapt LSLMs for long-speech inputs, it introduces a dynamic compression training approach, which exposes the model to short-speech sequences at varying compression ratios, thereby transferring the capabilities of LSLMs to long-speech tasks. To assess the long-speech capabilities of LSLMs, we develop a long-speech understanding benchmark called LongSpeech-Eval. Experiments show that our method exhibits strong performance in both long-speech and short-speech tasks, while greatly improving inference efficiency  [^1] The Code is at https://github.com/ictnlp/FastLongSpeech.git..
@@ 1 Introduction
To handle speech inputs, traditional methods [8, 9] typically employ a cascaded pipeline, where speech is first transcribed into text and then processed by LLMs. However, these approaches suffer from error propagation and discard valuable paralinguistic information [10]. To overcome these drawbacks, recent research [11, 12, 13] has shifted towards an end-to-end paradigm, enabling LSLMs to directly process and reason with speech signals. These methods can be broadly divided into two categories. On one hand, some approaches [4, 14, 15, 16] such as Qwen2-Audio [5] align the output spaces of pre-trained audio encoders with the embedding of LLMs, which allows speech inputs to be accommodated while transferring partial capabilities of the employed LLMs. At the same time, other methods [11, 17] involves discretizing the speech to discrete units, which allows LSLMs to handle speech units similar to text tokens. Despite these advancements, current LSLMs are largely constrained to processing short speech segments, typically under 30 seconds [5, 18]. Only a few LSLMs [12] have achieved a processing duration of 30 minutes on speech summarization tasks by relying on the construction of extensive, specialized training datasets.
@@ 2 Background
Connectionist Temporal Classification (CTC) [23] is widely used in Automatic Speech Recognition (ASR), where the input is typically longer than the corresponding output [26]. To align speech with the transcript, CTC introduces a blank token into the vocabulary and defines the possible output as an alignment $\mathbf{a}$. Each time step in this alignment corresponds to either a blank token or a non-blank token. The alignment has the same length as the speech sequence and can be reduced to the final transcript through a collapse function $\Gamma^{-1}$. During training, CTC employs an efficient dynamic programming algorithm to maximize the probability of all possible alignments corresponding to the ground-truth transcript $\mathbf{c}$:
@@ 3.2 Iterative Fusion
r(m)=T(m)T(m+1).r(m)=T(m)-T(m\!+\!1). (6)
Subsequently, we utilize the similarity metric between adjacent frames, as defined in Eq.(4) to identify the $r(m)$ most similar pairs of adjacent frames. The consecutive identified frames are grouped into a span, as illustrated in Figure 1. For each span, we employ a weighted fusion approach, leveraging the content density in Eq.(3) as weights to merge all frames within the span into a single compressed speech frame. This process yields a new sequence of speech representations with length T(m+1)T(m\!+\!1), comprising both the newly condensed frames and the remaining uncompressed frames. If T(m+1)T(m\!+\!1) still exceeds the target length $L$, we initiate another iteration. Otherwise, the resulting speech representations from this round constitute the final condensed representations $\mathbf{h}^{\prime}$.
@@ 3.3 Training Method
In the second stage, we introduce a novel dynamic compression training method. The introduction of this method is based on two considerations. First, it enables the LLM to adapt to condensed representations $\mathbf{h}^{\prime}$ with different compression ratios. Second, the current $\mathbf{s}$, $\mathbf{x}$, $\mathbf{y}$> triplet training data primarily contains short-speech clips, which are typically under 30 seconds in duration [29]. By sampling the length $L$ of condensed representations [30], LSLM can maintain its perception of speech sequences corresponding to the length of its speech window, without avoiding excessive bias towards overly condensed speech sequences. The dynamic compression training method is as follows:
@@ 4.1 Datasets
Speech Information Retrieval: We introduce SPIRAL-H benchmark, which is introduced in the paper of SpeechPrune [42] for long speech information retrieval.
@@ 4.2 System Settings
Cascaded first uses Whisper-Large-V3 [22] to transcribe the audio, and then pass the resulting context text to Qwen-7B-Chat (since Qwen2-Audio-7B is based on Qwen-7B-Chat).
@@
(a) Performance of diverse speech fusion methods in the short-speech spoken QA tasks. The score is derived from the LLM evaluating the quality of responses based on the questions and ground-truth answers. The baseline model utilizes a speech window of 750 frames. For the methods other than Baseline, we regulate the compression ratio by adjusting the target length $L$ of the condensed speech representations. In the experiments, a smaller value of $L$ corresponds to a higher compression ratio. A higher score indicates a better quality of the responses.
@@ 4.3 Main Results
For the short-speech spoken QA task, Figure 4(a) illustrates the performance of various methods across three datasets. Our FastLongSpeech method consistently outperforms other methods on all three datasets under different speech compression ratios, maintaining high response quality even at a 30-fold compression ratio ($L$ = 25). Unlike the Random method, which arbitrarily discards speech representations, other methods consider all temporal information when compressing speech representation, resulting in improved generation quality [29]. Compared to AvgPool and MostSim methods, our method more effectively eliminates redundant information while preserving highly informative speech representations, leading to better performance across various compression ratios. Notably, when $L$ equals 12, other speech fusion methods exhibit similarly suboptimal performance, while our method maintains a substantial performance advantage. We attribute this superiority to our novel iterative fusion strategy and dynamic compression training approach. Furthermore, compared to vanilla Qwen2-Audio, our method achieves comparable performance with a shorter sequence of speech representations, demonstrating higher efficiency.
@@ 5.2 Inference Efficiency
Ours (LL=400) 3.80 8.54
Ours (LL=200) 3.87 5.64
Ours (LL=100) 3.71 4.17
@@
In short-speech tasks, our approach demonstrates performance comparable to vanilla Qwen2-Audio while requiring only half the computational resources. When the allocated computational resources are increased, we can achieve better results. Notably, the computational costs decrease as the compression ratio increases. This not only demonstrates the better efficiency of our model but also highlights its ability to balance generation quality and inference efficiency.
@@
The advantages of our method become even more pronounced in long-speech tasks, where our method achieves better generation quality than NTK-ROPE, with a 70% reduction in runtime and a 60% decrease in computational costs. Compared to the cascaded method, it even achieves a speedup of more than sevenfold, underscoring its substantial efficiency advantage for processing long-form speech. This further shows the effectiveness of our method in handling long-speech inputs. For spoken dialogue understanding, emotion recognition and speech information retrieval tasks, please refer to the Appendix E and F.
@@ 5.3 Content of Condensed Representations
Ours (LL=750) 4.04 7.02 11.76
Ours (LL=400) 4.08 7.17 11.77
Ours (LL=200) 4.36 7.40 12.70
Ours (LL=100) 27.12 24.61 23.69
Beyond the spoken QA, spoken dialogue understanding and emotion recognition tasks, we extend our evaluation to the ASR task, which requires precise transcription of the entire speech content [47]. Through this task, we explore variations in condensed representations across different compression ratios. Table 5 demonstrates the ASR performance of Qwen2-Audio and our method. At low compression ratios (LL=400), FastLongSpeech performs comparably to Qwen2-Audio, demonstrating the effectiveness of our dynamic compression training and iterative fusion strategy in preserving speech content. Unlike Qwen2-Audio, our method does not require substantial post-processing to extract the transcript, with strong instruction following abilities. At higher compression ratios (LL=100), FastLongSpeech slightly trails Qwen2-Audio in ASR but maintains comparable results in spoken QA, as illustrated in Figure 4(a). This indicates that while our approach demonstrates applicability across diverse tasks, the optimal compression ratio is inherently task-dependent. Therefore, achieving an effective balance between efficiency and effectiveness thus necessitates careful calibration and a thorough assessment of resource constraints.
@@ 6 Related Work
Long sequence modeling presents challenges across diverse domains, including text, video, and speech. The approaches to long-context modeling vary depending on the type of the inputs. For extended text sequences, researchers explored methods such as position interpolation and extrapolation [51], sliding window [52], continuous fine-tuning on long-text data [53], and native sparse attention [54]. To address long-video processing, recent works leverage frame selection or merging strategies [55], as well as vision token merging techniques [56]. In the realm of speech processing, early methods focus on enhancing the performance of ASR [57] and speech translation [58] through speech compression techniques. More recently, FastAdaSP [40] mitigates inference overhead by performing token selection within LLMs. Concurrently, Speechprune [29] employs a token selection strategy to extend the effective speech window of Qwen2-Audio to 90 seconds for Speech Information Retrieval task. StreamUni [42] achieves real-time speech translation for long speech streams by integrating a segmentation strategy and a policy-decision module.
@@ 7 Conclusion
Acknowledgement
We gratefully acknowledge all the reviewers for their valuable comments and suggestions. This work was supported by the Natural Science Foundation of Beijing, China (Grant No. L257006).
@@ Appendix E Applicability of Our Method to Vanilla LSLMs
Ours (LL=400) 4.13 8.25
Ours (LL=200) 3.92 5.46
Ours (LL=100) 3.62 4.06
Ours (LL=50) 3.16 3.35
To demonstrate the effectiveness and robustness of our method, we first extend our experiments to spoken dialogue understanding task. For this task, we conduct experiments on vanilla Qwen2-Audio using speech_dialogue_QA_fisher of AIR-Bench [38]. As shown in Table 7, our method effectively balances performance and inference efficiency. Notably, at lower compression ratios ($L$ = 200), our approach demonstrate comparable performance to the vanilla Qwen2-Audio model with a 50% reduction in computational costs. Moreover, even at a higher compression ratio of 15x (LL=50), our method still maintains robust performance. These findings underscore the efficacy and versatility of our iterative fusion strategy.
@@ Appendix F Applicability of Our Method to Other Tasks
Table 10: The experiments on the SPIRAL-H dataset. The performance is measured with accuracy metric.
Method Prune Rate (%)($\uparrow$) Accuracy (%) ($\uparrow$)
SpeechPrune 60.00 63.77
FastLongSpeech ($L=750$) 65.88 76.81
Beyond emotion recognition, we also conduct additional experiments on the SPIRAL-H [^11] https://github.com/linyueqian/SPIRAL_Dataset dataset, a benchmark designed for long speech information retrieval. On this dataset, we follow the experimental setup [29] and compare model performance under similar speech embedding pruning rates. As shown in the table 10, our method achieves better performance with fewer speech embeddings, demonstrating its superiority and efficiency in modeling long speech inputs.
Table 11: The performance of FastLongSpeech with varying context lengths.
$\mathbf{L}$ 200 400 750 1200 4000
Score ($\uparrow$) 2.47 2.90 3.55 3.66 3.59
Appendix G Extending Maximum Context Length
We also explore the performance of FastLongSpeech on the LongSpeech-Eval dataset with varying context lengths $\mathbf{L}$. The results are shown in Table 11. When $L$ is less than 750, the model exhibits increasing performance with longer context, as it is trained under dynamically varying compression ratios in this range. When $L$ equals 1200, although the model is not explicitly trained for this length, it still achieves strong performance, indicating good generalization beyond the training regime. When increases to 4000, performance slightly declines. This is expected, as the model is not exposed to such long contexts during training, despite having a larger speech context window. We think our FastLongSpeech can achieve better performance with longer effective context length as the long-speech training data becomes more available.